Grade upto college level Physical Chemistry

What is the molarity and molarity of a 13% solution (by weight) of suphuric acid with a density of 1.02 g/ml? To what volume should 100 ml of this acid be diluted in order to prepare a 1.5 N solution?

Shane Macguire
12 Years agoGrade upto college level

1 Answer

Deepak Patra
12 Years ago
Hello Student,
Please find the answer to your question
(i) Molarity = Moles of solute/Volume of solution in L
(ii) N1V1 = N2V2
A 13% solution (by weight) contains 13 g of solute (i.e H2SO4) per 100 gm of solution
Moles of solute = Mass of H2SO4/M. wt. of H2SO4 = 13/98 = 0.1326
Volume of solution in L
= Mass of solution/density of solution *1000 = 100/1.02 * 1000 = 0.0980 Litre
∴ Molarity of solution = 0.1326/0.0980 = 1.35 M
Again, Molality = Moles of solute/Mass of solvent in kg
Mass of solute in 100 ml of solution = 13 g [13% solution]
Mass of solvent = Mass of solution – Mass of solvent
= 100 – 13 = 87 g
∴ Molality = 13/98/87/100 = 1.57 m
Normality = Miolarity * Mol. wt/Eq. wt. or 1.35 * 98/49 = 2.70 N
N1 = 2.70, V1 = 100 ml, N2 = 1.5, V2= ? [∵ Eq. wt = 98/2 2H2SO4 = 49]
N1V1 = N2V2
2.70 * 100 = 1.5 * v2­
Or V2 = 2.70 *100/1.5 = 180 ml.
∴ 100 ml of this acid should be diluted to 180 ml to prepare 1.5 solution.

Thanks
Deepak patra
askIITians Faculty